# Revisiting Experience Replayable Conditions

### v1

@@ Appendix A Details of implementation
The policy and value functions are approximated by a fully-connected neural networks consisting of two hidden layers with 100 neurons for each. As activation functions, Squish function [1, 28] and RMSNorm [56] are combined. AdaTerm [18], which is the noise-robust optimizer, is adopted for robustly optimizing parameters against noises caused by bootstrapped learning in RL. Similarly, target networks that are updated by CAT-soft update [22] are employed to stabilize learning (and to prevent learning speed degradation). In A2C, the target networks is applied to both the policy $\pi$ and the value function $V$, while in SAC they are applied only to the action value function $Q$. In addition, A2C enhances output continuity by using L2C2 [23] with default parameters, although SAC does not so due to reproduction of its standard implementation.

### v2

@@ Acknowledgments
Declarations
[…]
Competing Interests
[…]
Author Contributions
Taisuke Kobayashi contributed to everything for this paper: Conceptualization, Methodology, Software, Validation, Investigation, Visualization, Funding acquisition, and Writing.
[…]
Compliance with Ethical Standards
The data used in this study was exclusively generated by the author. No research involving human participants or animals has been performed.
@@ Appendix A Details of implementation
The policy and value functions are approximated by the fully-connected neural networks consisting of two hidden layers with 100 neurons for each. As activation functions, Squish function [2, 31] and RMSNorm [63] are combined. AdaTerm [21], which is the noise-robust optimizer, is adopted for robustly optimizing parameters against noises caused by bootstrapped learning in RL. Similarly, the target networks that are updated by CAT-soft update [30] are employed to stabilize learning (and to prevent learning speed degradation). In A2C, the target networks are applied to both the policy $\pi$ and the value function $V$, while in SAC they are applied only to the action value function $Q$. In addition, A2C enhances output continuity by using L2C2 [25] with default parameters, although SAC does not so due to reproduction of its standard implementation.
@@
Appendix B Application to PPO
(a) Quadrotor
(b) Fixedwing
(c) Returns after learning
Figure 9: Returns with PPO on PyFlyt: From (a) and (b), it can be found that the mining trick improves learning speed, while the counteraction trick facilitates exploration, which yielded the maximum and stable return on Fixedwing in (c).
The proposed stabilization tricks are applied to PPO [47], a latest on-policy algorithm other than A2C. PPO multiplies the policy improvement by the policy likelihood ratio using importance sampling, and it is clipped to force the gradient to zero if it is excessive. The recommended clipping threshold, $0.2$, is employed. Since this paper uses an ER that accumulates empirical data for the simplest single transition, GAE [46], which is often used in conjunction with PPO, is ignored. In addition, to check the regularization effects by the stabilization tricks, an extra regularization term, i.e. the policy entropy, is also omitted. As mentioned in Introduction, as PPO is empirically ER-applicable, it is possible to quantitatively compare the learning trends and changes in the final policies due to each stabilization trick.
With the above setup, the results of solving QuadX-Waypoints-v2 and Fixedwing-Waypoints-v2 provided by PyFlyt [52] are summarized below. Note that these tasks intend to control different drones (i.e. a quadrotor and a fixed-wing drone), respectively. The learning curves for 7 trials and test results after learning in the four conditions diverging with and without the two stabilization tricks are depicted in Fig. 9.
First, the vanilla PPO shows the remarkable but expected results. Although PPO is originally considered an on-policy algorithm, as described in the text, it achieved learning of both tasks in combination with ER even without the addition of the proposed stabilization tricks. This is because, as discussed in Section 6.1, PPO performs regularization and clipping heuristically such that $\pi\simeq b$ (i.e. on-policyness) holds. In fact, PPO with only one of the stabilization trick did not saturate the corresponding internal parameter, indicating that ERC was satisfied without the lack of its functionality. Note, however, that PPO by itself may not be sufficient to satisfy ERC, since GAE, which was omitted from the implementation this time, relies more heavily on the behavior policy than the simple advantage function (i.e. TD error).
Next, the contributions of the proposed stabilization tricks are confirmed. First, it is easy to see that the mining trick increases learning speed, while the counteraction trick tends to decrease it. This may be due to the fact that the direction of policy improvements becomes clearer by making the mining trick replay only the empirical data that are useful for learning, while the counteraction trick restricts policy improvements by binding $\pi\simeq b$. For the former, actually, the scale of TD error, which implies the learning direction, was increased when the mining trick was added.
On the other hand, the counteraction trick seems to improve the exploration capability in exchange for the learning speed. The return at the end of learning was maximized by including the counteraction trick on Fixedwing, although the return on Quadrotor with it was lower than others due to slow convergence. This may be due to the increase in entropy of $\pi$ by regularizing it to various the behavior policies in the replay buffer. In fact, the addition of the counteraction trick yielded the decrease of $\ln\pi$, which corresponds to the negative entropy, during learning.